# Antipodality of Spherical Designs with Odd Harmonic Indices

### v5

@@ 3. Interval Designs
Remark $3.$
Although in this paper we restrict ourselves to interval designs for the normalized Lebesgue measure on [1,1][-1,1], the notion of an interval design naturally extends to an arbitrary measure on an interval. In this broader setting, a challenging problem is the construction of interval designs with rational points and weights for the ultraspherical measure (see Bannai et al. [4, p. 208, Problem 2]). It was Mishima et al. [13] who first observed Proposition 3.2 for $m\leq 3$ and thereby established, together with explicit constructions, a classification theorem of rational 55-designs with 66 points for the Chebyshev measure $(1-t^{2})^{-1/2}dt/\pi$.
In Proposition 3.2, the set $\{x_{1},\dots,x_{n}\}$ is regarded as a multiset. This is equivalent to consider a function $\lambda$ on $\{x_{1},\dots,x_{n}\}$ taking values in the set of positive integers. If we allow $\lambda$ to be a real-valued function, we can also conclude $f$ to be an even function under a stronger hypothesis.
@@ 5. Optimality
Let $f(x)$ be the monic polynomial of degree $2m$ having AA as its set of roots. We can choose a small $\varepsilon>0$ in such a way that the polynomial $g(x)=f(x)+\varepsilon$ has $2m$ roots in $(-1+\frac{1}{2m},1-\frac{1}{2m})\setminus\{\frac{1}{2m}\}$. Let $A^{\prime}$ be the set of roots of $g(x)$. Then, by the relation between roots and coefficients,
@@
Since AA and $(A-\tfrac{1}{2m})\cup\{1\}$ are both symmetric, $p_{\mathrm{odd}}(A)=0$, and
@@
Since $A^{\prime}$ is also symmetric, we can reverse the above calculation with AA replaced by $A^{\prime}$, and we conclude
@@
The second assertion then follows from Theorem $4.$ ∎
@@
Although Proposition 5.6 shows that the inequality $n\leq 2m$ in the assumption of Theorem 4.4 cannot be relaxed, it does not assert the existence of a non-antipodal spherical $T_{m}$-design with $n$ points for even $n>2m$. We show that Proposition 5.6 cannot be strengthened to include the case where $n$ is even, by proving the non-existence of a non-antipodal $T_{2}$-design with 66 points on $\mathbb{S}^{1}$. Note that Proposition 5.6 does imply the existence of a non-antipodal $T_{m}$-design with $n\geq 4m+2$ points as a union of two disjoint $(2m+1)$-gons.
@@
There is no non-antipodal $T_{2}$-design on $\mathbb{S}^{1}$ of the cardinality 66.
@@
Let $X\subset\mathbb{S}^{1}$ be a spherical $T_{2}$-design with $|X|=6$ points. We will show that $X$ is antipodal. Suppose not. Then $X\cap(-X)=\emptyset$ by Theorem $4.$ In particular, we may assume by rotating that $(\pm x_{1},y_{1})\in X$ for some $x_{1}$ with $0<x_{1}<1$, $0<y_{1}$. The projection $\bar{X}$ of $X$ onto the $x$-coordinate gives a weighted interval $T_{2}$-design, and by removing $\pm x_{1}$ with multiplicity 11, we obtain a weighted interval $T_{2}$-design $X^{\prime}$ consisting of 44 points. By Proposition 3.2, these 44 points form a symmetric multiset $X^{\prime}$. Suppose first that 44 points of $X^{\prime}$ are not distinct. If there are only 33 distinct points, then one of the points is 0 with multiplicity 22, which implies $(0,\pm 1)\in X$. This contradicts the fact that $X\cap(-X)=\emptyset$. If there are only two distinct points, then we have $X^{\prime}=\{\pm x_{2}\}$, which also contradicts the fact that $X\cap(-X)=\emptyset$. We next suppose that $X^{\prime}$ consists of 44 distinct points $\{\pm x_{2},\pm x_{3}\}$. Then we may assume

### v6

@@ 3. Interval Designs
Remark 3.3.
Although in this paper we restrict ourselves to interval designs for the normalized Lebesgue measure on [1,1][-1,1], the notion of an interval design naturally extends to an arbitrary measure on an interval. In this broader setting, a challenging problem is the construction of interval designs with rational points and weights for the ultraspherical measure (see Bannai et al. [4, p. 208, Problem 2]). It was Mishima et al. [13] who first observed Proposition 3.2 for $m\leq 3$ and thereby established, together with explicit constructions, a classification theorem of rational $5$-designs with $6$ points for the Chebyshev measure $(1-t^{2})^{-1/2}dt/\pi$.
In Proposition 3.2, the set $\{x_{1},\dots,x_{n}\}$ is regarded as a multiset. This is equivalent to consider a function $\lambda$ on $\{x_{1},\dots,x_{n}\}$ taking values in the set of positive integers. If we allow $\lambda$ to be a real-valued function, we can also conclude $\lambda$ to be an even function under a stronger hypothesis.
@@ 5. Optimality
Let $f(x)$ be the monic polynomial of degree $2m$ having $A$ as its set of roots. We can choose a small $\varepsilon>0$ in such a way that the polynomial $g(x)=f(x)+\varepsilon$ has $2m$ roots in $(-1+\frac{1}{2m},1-\frac{1}{2m})\setminus\{\frac{1}{2m}\}$. Let $A^{\prime}$ be the set of roots of $g(x)$. Then, by the relation between roots and coefficients,
@@
Since $A$ and $(A-\tfrac{1}{2m})\cup\{1\}$ are both symmetric, $p_{\mathrm{odd}}(A)=0$, and
@@
Since $A^{\prime}$ is also symmetric, we can reverse the above calculation with $A$ replaced by $A^{\prime}$, and we conclude
@@
The second assertion then follows from Theorem 4.4. ∎
@@
Although Proposition 5.6 shows that the inequality $n\leq 2m$ in the assumption of Theorem 4.4 cannot be relaxed, it does not assert the existence of a non-antipodal spherical $T_{m}$-design with $n$ points for even $n>2m$. We show that Proposition 5.6 cannot be strengthened to include the case where $n$ is even, by proving the non-existence of a non-antipodal $T_{2}$-design with $6$ points on $\mathbb{S}^{1}$. Note that Proposition 5.6 does imply the existence of a non-antipodal $T_{m}$-design with $n\geq 4m+2$ points as a union of two disjoint $(2m+1)$-gons.
@@
There is no non-antipodal $T_{2}$-design on $\mathbb{S}^{1}$ of the cardinality $6$.
@@
Let $X\subset\mathbb{S}^{1}$ be a spherical $T_{2}$-design with $|X|=6$ points. We will show that $X$ is antipodal. Suppose not. Then $X\cap(-X)=\emptyset$ by Theorem 4.4. In particular, we may assume by rotating that $(\pm x_{1},y_{1})\in X$ for some $x_{1}$ with $0<x_{1}<1$, $0<y_{1}$. The projection $\bar{X}$ of $X$ onto the $x$-coordinate gives a weighted interval $T_{2}$-design, and by removing $\pm x_{1}$ with multiplicity $1$, we obtain a weighted interval $T_{2}$-design $X^{\prime}$ consisting of $4$ points. By Proposition 3.2, these $4$ points form a symmetric multiset $X^{\prime}$. Suppose first that $4$ points of $X^{\prime}$ are not distinct. If there are only $3$ distinct points, then one of the points is 0 with multiplicity $2$, which implies $(0,\pm 1)\in X$. This contradicts the fact that $X\cap(-X)=\emptyset$. If there are only two distinct points, then we have $X^{\prime}=\{\pm x_{2}\}$, which also contradicts the fact that $X\cap(-X)=\emptyset$. We next suppose that $X^{\prime}$ consists of $4$ distinct points $\{\pm x_{2},\pm x_{3}\}$. Then we may assume